OpenStudy (anonymous):

-3|x-2|+10=12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No Solution

OpenStudy (anonymous):

@AravindG HELP ME PLEASE STEP BY STEP

OpenStudy (anonymous):

\[-3|x-2|=2\] now we square this.. the only way to get rid of the absoute sign is to square it

OpenStudy (anonymous):

or make two equations out of it.

OpenStudy (anonymous):

ok wait spuare the hole thing

OpenStudy (anonymous):

yes see the above simplified orm

OpenStudy (anonymous):

\[\left[-3|x-2|\right]^2=2^2\]

OpenStudy (anonymous):

ok wat about the + 10

OpenStudy (anonymous):

I subtracted it from the two sides

OpenStudy (anonymous):

notice that the right side became 12-10=2

OpenStudy (anonymous):

where did u get the 12 frm

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

where did u get the 12 from

OpenStudy (anonymous):

\[-3|x-2|+10=12\qquad\text{....given}\\ \qquad\qquad-10\quad-10 \]

OpenStudy (anonymous):

ok srry im just learning this ok and then gives u this [−3|x−2|]^2=2^2

OpenStudy (anonymous):

that is ok. well, this first gives \[-3|x-2|=2\] so, we the square it like you just wrote above

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(-3)^2(x-2)^2=(2^2)\] notice that the absolute sign has disappeared

OpenStudy (anonymous):

so 9*2x-4=4

OpenStudy (anonymous):

or to avoid squaring, you can remove the abolute sign in two ways...... \[-3|x-2|=2\implies|x-2|=-{3\over2}\]

OpenStudy (anonymous):

how did u get 3/2

OpenStudy (anonymous):

my bad... \[\Large{|x-2|=\color{red}{-}{2\over3}}\]

OpenStudy (anonymous):

ooh its ok but how did u get that

OpenStudy (anonymous):

divide the two sides by "3"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[{-3|x-2|\over-3}={2\over-3}\]

OpenStudy (anonymous):

this gives the above form.. now, by definition of ABSOLUTE value, can it ever become negative?

OpenStudy (anonymous):

HEY STOP STOP STOP I WROTE THE PROBLEM DWN WRONG SRRY THIS IS THE PROBLEM −3|2x + 6| = −12

OpenStudy (anonymous):

ok... you mean no "10" in there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

this is so much different than what you first asked for

OpenStudy (anonymous):

ok.. proceeding similarly, get the absolute quantity on its own...

OpenStudy (anonymous):

ik b/c its 2 differnt problems in one

OpenStudy (anonymous):

\[\Large{-3|2x-6|=-12\\\text{divide both sides by -3}\\ |2x-6|=4\\\text{to get rid of absolute sign, we get two possible answers}\\\;\\ 2x-6=4\qquad{\rm AND}\qquad2x-6=-4\\ 2x=4+6\qquad\qquad\qquad2x=-4+6\\ 2x=10\qquad\qquad\qquad\quad2x=2\\ \boxed{x=5}\qquad\qquad\qquad\quad\boxed{x=1}} \]

OpenStudy (anonymous):

check, plug in "x=" those numbers in the given problem on the left side. That should give you the right side.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

did you understand?

OpenStudy (anonymous):

yes